A factory produces apple juice contained in a bottle of 1.5L. However, due to random fluctuations in the automatic bottling machine, the actual volume per bottle varies according to a normal distribution. It is observed that 10% of bottles are under 1.45L whereas 5% contain more than 1.55L. Calculate the mean and standard deviation of the volume distribution。


1
Expert's answer
2021-11-08T20:41:07-0500

Let "X=" the volume per bottle: "X\sim N(\mu, \sigma^2)"


"P(X<1.45)=P(Z<\dfrac{1.45-\mu}{\sigma})=0.1"

"\dfrac{1.45-\mu}{\sigma}\approx-1.28155"


"P(X>1.55)=1-P(Z\leq\dfrac{1.55-\mu}{\sigma})=0.05"

"P(Z\leq\dfrac{1.55-\mu}{\sigma})=0.95"

"\dfrac{1.55-\mu}{\sigma}\approx1.64485"

"\dfrac{1.45-\mu}{1.55-\mu}=\dfrac{-1.28155}{1.64485}"

"2.3850325-1.64485\mu=-1.9864025+1.28155\mu"

"2.9264\mu=4.371435"

"\mu=1.493793"

"\sigma=\dfrac{1.45-1.493793}{-1.2815}"

"\sigma=0.034173"

"mean=\mu=1.4938L"

"\sigma=0.0342L"